OpenStudy (lukecrayonz):

x=e^(3y) I got to ln(x)=3yln(e), and when I use an online calculator to help me, they just switch ln to log for no reason :'(

OpenStudy (amistre64):

wolfram notates ln as "log" unless stated otherwise

OpenStudy (lukecrayonz):

Ooooh =[

OpenStudy (amistre64):

what is it your trying to do with this?

OpenStudy (lukecrayonz):

Well i'm given y=e^3x, and im finding the inverse.

OpenStudy (lukecrayonz):

Solving for y by inputting x for y, y for x. x=e^3y, then solving. ln(x)=3yln(e)

OpenStudy (lukecrayonz):

Not sure what to do from there!

OpenStudy (amistre64):

ln and e are inverses of each other; so they cancel each other out, or rather they equate to "1" like 5/5: ln(e) = 1 e^(ln(1))=1

OpenStudy (amistre64):

once you know that, divide off that 3 and your set

OpenStudy (lukecrayonz):

So wait, just ln(x)/3=y?

OpenStudy (amistre64):

yep

OpenStudy (lukecrayonz):

Thanks :D

OpenStudy (amistre64):

youre welcome

OpenStudy (lukecrayonz):

I have another one D: I always just get stuck at the beginning.

OpenStudy (lukecrayonz):

x=log2y, is it 10^x/2=y?